40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Feb 5 21:42:28 2009, Yoichi, Update, PSL, My thoughts on ISS INMONPD_Spectrum_1-10Hz.pdf
    Reply  Fri Feb 6 09:52:35 2009, Kakeru, Update, PSL, Current shunt transfar function TF_CS_gain.pngTF_CS_phase.png
       Reply  Fri Feb 6 23:23:48 2009, Kakeru, Yoichi, Update, PSL, ISS is fixed Current_ShuntTF_gain.pngCurrent_ShuntTF_phase.png
          Reply  Mon Feb 9 16:02:42 2009, Yoichi, Update, PSL, PSL relative intensity noise RIN-13dB.pngRIN-21dB.png
             Reply  Tue Feb 10 23:36:25 2009, Kakeru, Update, PSL, PA current and laser output PA_current_output.png
                Reply  Wed Feb 11 07:28:25 2009, Yoichi, Update, PSL, PA current and laser output 
                   Reply  Wed Feb 11 23:51:53 2009, Kakeru, Update, PSL, PA current and laser output PA_current_output.png
Message ID: 1295     Entry time: Wed Feb 11 23:51:53 2009     In reply to: 1291
Author: Kakeru 
Type: Update 
Category: PSL 
Subject: PA current and laser output 
I attached a plot of ISS monitor PD and MOPA output to PA current.
The both end of PA current (26.0353[A] and 28.4144[A]) correspond to the slider value of -2.0 and 1.0 .
It looks that we must use MOPA with PA current below 27.5[A].
Attachment 1: PA_current_output.png  12 kB  | Hide | Hide all
PA_current_output.png
ELOG V3.1.3-